# Matteo Escudé

### v5

@@ Abstract
I study a model of information acquisition and transmission in which the sender’s ability to misreport her findings is limited. In equilibrium, the sender only influences the receiver by choosing to remain selectively ignorant, rather than by deceiving her about the discoveries. Although deception does not occur, I highlight how deception possibilities determine what information the sender chooses to acquire and transmit. I then turn to comparative statics, characterizing in which sense the sender benefits from her claims being more verifiable, showing this is akin to increasing her commitment power. Finally, I characterize sender- and receiver-optimal falsification environments.
@@ 1 Introduction
To inform their choices, individuals typically rely on the advice of experts with access to superior information. In the presence of conflicting interests, the latter will attempt to influence the former’s decisions in two ways: by acquiring information selectively and—as much as it is feasible—by misrepresenting their findings. This paper studies the interplay between these two modes of influence, exploring how the expert’s ability to misrepresent her discoveries determines what information she acquires and, ultimately, transmits to the receiver.
@@
These settings share three key features, which are also the essential components of the model I develop. First, there is a substantial conflict of interest between the ‘sender’ and the ‘receiver’. Because of this conflict, the former will use her ability to control information production and transmission to influence the latter’s decision in her favor, making her—in turn—wary of the advice.
@@
Third, the sender faces constraints on the extent to which she can misrepresent her findings. This may be because some of her claims are verifiable by the receiver or by a trusted third-party. She might, also, have reputational concerns, fearing that reports excessively different from what she finds may be ultimately discovered, damaging her reputation and future credibility. Additionally, she might be bound by disclosure regulation or contractual obligations. [^1] For example, reputational concerns are the key (albeit imperfect) reason behind the credibility of credit-rating agencies (see Mathis et al. (2009) for a discussion). Disclosure regulation plays a crucial role in financial reporting (see Leuz and Wysocki (2016) for a survey).
My contribution is twofold. First, I develop a novel model of information acquisition and transmission incorporating these three ingredients. I do so by marrying the verifiable-disclosure literature’s flexible approach to constrained communication with Bayesian persuasion’s belief-based methods.
Second, I show that the combination proves fruitful: I obtain a robust unraveling result across all possible misreporting constraints in a large (and natural) class and highlight how they determine what information the sender chooses to acquire. This, in turn, allows to address comparative statics questions, delineating the connection between more verifiability of the sender’s reporting with more commitment power in Bayesian persuasion problems with limited sender commitment. It also allows to characterize optimal verifiability environments: for the sender, this is a characterization of what constraints allow her to obtain the Bayesian persuasion outcome; for the receiver, they describe constraints that incentivise the sender to acquire (and transmit) full information in equilibrium.
Model outline.
Two players, a sender and a receiver, are initially equally uninformed about a binary (high or low) state. The sender can covertly and freely acquire information about the state and subsequently communicate with the receiver. The players have conflicting interests: the sender (weakly) benefits from inducing higher beliefs about the state in the receiver.
A key primitive of the model is the verifiability structure, which describes what the sender can say to the receiver following any given outcome of her information acquisition. Formally, if the sender discovers that the state is high with probability $s\in[0,1]$, she can send a message from the set $M(s)$ to the receiver.
Extreme verifiability structures help illustrate the environment. $M(s)=\{s\}$ describes a situation in which the sender is obliged to report her findings truthfully. This captures, for example, settings in which the outcome of the sender’s information acquisition is observed directly by the receiver or, alternatively, extremely stringent disclosure requirements. Conversely, if $M(s)=\bar{M}$ for some given (perhaps very large) set $\bar{M}$, what the sender learns does not constrain at all what she can say to the receiver. This describes well settings in which ‘talk is cheap’, which may arise because the sender’s acquired information is entirely unverifiable.
Beyond these extremes, the class of verifiability structures I consider is flexible enough to encompass a variety of environments in which verifiability is full ($s\in M(s)$ for each $s\in[0,1]$, $s\notin M(s^{\prime})$ for $s^{\prime}\neq s$, as in Grossman and Hart (1980), Grossman (1981) and Milgrom (1981)) or partial ($M(s)\subseteq M(s^{\prime})$ for some $s,s^{\prime}\in[0,1]$, as introduced in Okuno-Fujiwara et al. (1990)).
@@
Observe that there are two natural ways of interpreting the set-up in the context of canonical models of information production and transmission. It is a sender-receiver game with (partially) verifiable information, in which the sender’s type distribution is determined by the sender’s equilibrium covert information acquisition, rather than fixed and common knowledge. It is also a model of Bayesian persuasion with limited commitment, in which the sender can acquire information freely and covertly, but can only commit to sending a message from the set $M(s)$ following realization $s$. In the canonical Bayesian persuasion setting, the sender’s information acquisition is public instead.
Overview of the results.
@@
The first set of results concerns environments in which $M$ allows the sender to prove news better than the prior (PNBP). PNBP holds if there exists a message that, even when interpreted most skeptically by the receiver, leads to a payoff for the sender strictly above what she would obtain at the prior. [^2] See Definition 3 for a formal statement. To illustrate: full-verifiability environments always satisfy PNBP provided that the sender is not already obtaining her highest possible payoff at the prior; cheap talk settings never satisfy PNBP. Partial verifiability settings may or may not satisfy PNBP.
If PNBP holds, an unraveling result obtains (Theorem 1 and Corollary 1): in any equilibrium the receiver interprets on-path messages in the least favorable way for the sender and learns exactly what the latter discovered about the state. That is, in equilibrium the sender never acquires information in a way that would lead her to pool types in the communication phase. Importantly, the result highlights that such unraveling prevails even when verifiability is only partial (provided PNBP holds), which would typically fail were the sender’s private information exogenous. [^3] Failures of unraveling due to partial verifiability with exogenous private information are highlighted, for example, in Okuno-Fujiwara et al. (1990) and Hagenbach et al. (2014).
The fact that, in equilibrium, all information acquired by the sender is revealed to the receiver does not mean that the sender learns (and reveals) the value of the state. Indeed, the sender is still able to influence the receiver’s decision in her favor by choosing what information to acquire. Corollary 2 characterizes the value of (covert) information acquisition in this setting: if the sender can prove news better than the prior, the sender’s equilibrium value is unique, and corresponds to a skepticism-adjusted version of the concave envelope of Kamenica and Gentzkow (2011).
@@
If the sender cannot prove news better than the prior, I show that there is always a sender-preferred equilibrium in which she acquires and transmits no information (Proposition 1). This result demonstrates how the conflict of interest makes information control worthless, absent some minimal amount of verifiability.
I then turn to comparative statics. I study in what sense the sender benefits from being in an environment exhibiting ‘more verifiability’. I remark (in Section 4.3) that intuitive shifts towards more verifiability may in fact hurt the sender ex ante, if PNPB does not hold. This is caused by the lack of observability of the sender’s information acquisition strategy, which implies that in a more constrained environment, the equilibrium meaning of some messages may worsen with more verifiability, leading to a lower expected payoff.
@@
However, when PNBP does hold, this cannot happen in equilibrium, as Theorem 1 and Corollary 1 imply that the receiver is maximally skeptical. In this case, I provide a pre-order on the set of verifiability structures which characterizes those shifts that are unequivocally beneficial for the sender (Theorems 2(a) and 2(b)). The necessary and sufficient condition is easy to describe: the set of sender types that are able to separate from all lower types, regardless of the receiver’s belief, must grow in the sense of set inclusion. This guarantees that any equilibrium receiver belief distribution inducible in the lower-ranked verifiability structure is also inducible in the higher-ranked one.
@@
The contribution of the comparative statics theorem is twofold. First, the pre-order allows to compare a rich variety of common verifiability environments. For example, according to the pre-order, cheap-talk is minimal, while full-verifiability structures are maximal. Interval-partition verifiability structures are ranked if and only if one is a refinement of the other. [^4] Interval-partition verifiability structures are formally defined in Example 4. They capture settings in which there are ‘grades’ or ‘ratings’ that certifiably pool together sets of types. The specialization of the comparative statics theorem to this setting is discussed in Example 6 and Remark 2. Adding a new message—thereby allowing the sender to credibly communicate her acquired information more finely—always leads to a higher-ranked structure. [^5] See Example 7 and Remark 2 for the details. Second, the result can be viewed as providing an appropriate notion for comparing Bayesian persuasion environments with different forms of sender commitment. In this context, it characterizes the appropriate notion of ‘increasing commitment power’ which is valuable for the sender.
@@
Theorems 2(a) and 2(b) lead also directly to a characterization of sender-optimal verifiability (Proposition 2). Such structures allow the sender to attain the full-commitment value in equilibrium. The result therefore highlights how the full commitment assumption in the canonical Bayesian persuasion setting can be relaxed, without affecting the sender’s equilibrium payoff. The necessary and sufficient condition states that all sender types must be able to separate from all lower ones by means of a single message, no matter what the receiver believes. Given the monotone conflict of interest, this guarantees that separation is possible for all types in equilibrium, allowing the sender replicate the Bayesian persuasion outcome.
@@
Finally, I characterize receiver-optimal structures (Proposition 3). These take a simple form: only the very best news (that is: claiming that the state is high with probability $1$) is considered credible, while all inconclusive news are fully unverifiable. Every structure of this form is receiver-optimal because it incentivizes (perfect) information acquisition on the part of the sender, which she then fully reveals in equilibrium.
@@ 1.1 Literature
A large literature, starting with the seminal work of Grossman and Hart (1980), Grossman (1981), Milgrom (1981) and Crawford and Sobel (1982), explores the role of verifiability (or lack thereof) in sender-receiver games. In this literature, the sender’s type distribution is fixed and common knowledge; in my case, the players start equally uninformed and the sender can privately acquire information before communicating with the receiver.
In the context of this literature with exogenous private information, several papers have pointed out that the classical unraveling results break down when information in only partially verifiable [^6] The classical unraveling argument was first provided by Grossman and Hart (1980), Grossman (1981) and Milgrom (1981). Failures of full information revelation due to partial verifiability are discussed, for example, in Okuno-Fujiwara et al. (1990), Lipman and Seppi (1995), Mathis (2008), Hagenbach et al. (2014) and Rappoport (2022). Unraveling may fail for other reasons, such as uncertainty about what the sender can disclose, as in the literature initiated by Dye (1985) and Jung and Kwon (1988) or disclosure costs, as in Jovanovic (1982) and Verrecchia (1983). (that is, when not all sender types have access to a message that separates them from all other types regardless of the receiver’s belief). In contrast, I provide conditions such that full revelation of acquired information obtains (Theorem 1 and Corollary 1) even in settings with partial verifiability, in which it would fail if the sender’s private information were exogenous, rather than acquired.
@@
A growing literature augments sender-receiver games by endogenizing the sender’s private information. It includes papers studying sender learning in cheap-talk models, [^7] For example, see Pei (2015) and Deimen and Szalay (2019); in the first information acquisition is covert (and costly), in the second it is overt. Earlier, Austen-Smith (1994) explored this question in a setting in which the sender either learns the value of the state perfectly, or not at all; in his setting, information acquisition is covert, but the sender can prove that she is informed. settings with full verifiability, [^8] An early example is Matthews and Postlewaite (1985), in which the sender faces a binary choice between a fully informative and verifiable signal and a fully uninformative and unverifiable signal. More recently, Gentzkow and Kamenica (2017) show that under full verifiability the sender perfectly reveals all acquired information. gradual evidence acquisition, [^9] For example, Argenziano et al. (2016), Felgenhauer and Loerke (2017), Janssen (2018), Herresthal (2022). or Dye (1985)-like environments. In this last category, my paper is related to Ben-Porath et al. (2018) and in particular to DeMarzo et al. (2019), as they also focus on how the lack of observability of the sender’s choice influences equilibrium outcomes. [^10] I further comment on the connection with their results in Section 3, after the statement of Theorem 1 and Corollary 1. Information acquisition with Dye evidence is also studied in Kartik et al. (2017) and Shishkin (2022).
In these papers, the only unverifiable message is ‘non-disclosure’, whose meaning the sender can worsen by making a different, unobservable, choice (of project, in the former paper, and of test, in the latter). In my model, the meaning of several messages (potentially of all of them) must to be jointly determined in equilibrium. [^11] That is: the receiver’s belief following any $m$ for which $M^{-1}(m)$ is not a singleton is determined in equilibrium. Another key difference is that in these papers the sender’s payoff is linear in the receiver’s belief. It follows that she is indifferent among all feasible distributions of the receiver’s posterior. The ‘selective information acquisition’ persuasion motive present in my model is therefore absent in these papers.
My complementary approach to the problem allows to speak directly to a variety of canonical settings121212 See Section 2.2 for a discussion of how these special cases fit into the general model. in the communication literature which cannot be captured by Dye (1985)-style evidence, pointing to a a novel property (PNBP, see Definition 3) which determines whether unraveling obtains (Theorem 1 and Corollary 1), if information control can at all be beneficial for the sender (Proposition 1) and—finally—whether ‘more verifiability’ favors or potentially harms the sender (Theorems 2(a) and 2(b), and Section 4.3).
@@
I also contribute to the literature exploring the consequences of relaxing the commitment assumption in Kamenica and Gentzkow (2011). Some papers allow the sender to alter negative results with some positive probability (for example, Lipnowski et al. (2022), Fréchette et al. (2022) and Min (2021)) while others study models of Bayesian persuasion with lying costs. In this latter category, Nguyen and Tan (2021) analyze a setting in which the sender’s information acquisition strategy is observable by the receiver while in Guo and Shmaya (2021)’s set-up, as in mine, it is not. Both provide conditions on the costs such that the sender-preferred equilibrium exhibits full information revelation.
@@
In contrast to these papers, in my model the sender faces a constraint on what she can or cannot communicate depending on what she discovers. This distinct formalization of partial commitment allows me to bridge the gap between the Bayesian persuasion set-up and models of verifiable disclosure. This, in turn, leads to a characterization of when ‘more verifiability’ of the sender’s acquired information it is akin to her having ‘more commitment power’.
@@
I introduce the model in the next section. The skepticism and unraveling result is provided in Section 3, alongside with a discussion of the value of persuasion with partial verifiability. Section 4 contains the comparative statics result, characterizing changes in verifiability that are beneficial for the sender, and the characterization of sender-optimal verifiability structures. Receiver-optimal structures are characterized in Section 5.

### v6

@@ Abstract
I study a model of information acquisition and transmission in which the sender’s ability to misreport her findings is limited. The sender learns covertly, so a key observation is that in equilibrium she must be deterred from undetectably worsening the meaning of the messages she sends. This force substantially disciplines equilibrium beliefs and behavior: the receiver is maximally skeptical of what the sender claims and learns all discovered information. I exploit these equilibrium properties to characterize the sense in which the sender benefits from her claims being more verifiable, showing that this is akin to increasing her commitment power. Finally, I identify sender- and receiver-optimal falsification environments.
@@ 1 Introduction
To inform their choices, individuals typically rely on the advice of experts with access to superior information. In the presence of conflicting interests, the latter will attempt to influence the former’s decisions in at least two ways: by acquiring information selectively and—as far as is feasible—by misrepresenting their findings. This paper studies the interplay between these two modes of influence, exploring how the expert’s ability to misrepresent her discoveries influences what information she acquires and, ultimately, transmits to the receiver.
@@
These settings share three key features, which are also the essential components of the model I study. First, there is a substantial conflict of interest between the ‘sender’ and the ‘receiver’. Because of this conflict, the sender will use her ability to control information production and transmission to influence the receiver’s decision in her favor, making the latter—in turn—wary of the advice.
@@
Third, the sender faces constraints on the extent to which she can misrepresent her findings. This may be because some of her claims are verifiable by the receiver or by a trusted third party. She might also have reputational concerns, fearing that reports excessively different from what she finds may be ultimately discovered, damaging her reputation and future credibility. Additionally, she might be bound by disclosure regulation or contractual obligations. [^1] For example, reputational concerns are the main (albeit imperfect) reason behind the credibility of credit-rating agencies (see Mathis et al. (2009) for a discussion). Disclosure regulation plays a crucial role in financial reporting (see Leuz and Wysocki (2016) for a survey).
@@
The model I develop incorporates these three ingredients by marrying the verifiable-disclosure literature’s flexible approach to constrained communication with Bayesian persuasion’s belief-based methods. My set-up can therefore be viewed from two different angles. It is a sender-receiver game with (partially) verifiable information, in which the sender’s type distribution is determined by her own equilibrium covert information acquisition, rather than fixed and common knowledge. It is also a model of Bayesian persuasion with limited commitment, in which the sender can acquire information freely and covertly, but can only commit to sending a message from a fixed set for each realization of her private signal.
Combining these two approaches proves fruitful: I obtain an unraveling result across all possible misreporting constraints in a large class, which highlights how these constraints shape what information the sender chooses to acquire. More specifically, since the sender learns covertly, she must be deterred in equilibrium from undetectably worsening the meaning of the messages she is expected to send. This serves as a key disciplining force, ensuring that the sender seeks only information that she can then credibly transmit, given the misreporting constraints.
This insight, in turn, allows me to address comparative statics questions, delineating the connection between more verifiability of the sender’s reporting and more commitment power in Bayesian persuasion problems with limited sender commitment. It also delivers a characterization of optimal verifiability environments: for the sender, this is a characterization of the constraints that allow her to obtain the Bayesian persuasion outcome; for the receiver, they describe constraints that incentivize the sender to acquire (and transmit) full information in equilibrium.
1.1 Example and overview of the results
Two players, a sender and a receiver, are initially uninformed about a binary state, taking values in $\{0,1\}$. They share a common prior belief $\overline{p}=1/4$ that the state is $1$. The sender can covertly and freely acquire information about the state by privately choosing a signal—that is: a mean-$1/4$ distribution of posterior beliefs—and observing its realization.
An important primitive of the model is the verifiability structure, which describes what the sender can say to the receiver following any given outcome of her information acquisition. Formally, if the sender discovers that the state is $1$ with probability $s\in[0,1]$, she can send a message from the set $M(s)$ to the receiver. [^2] The fact that the sender’s posterior belief uniquely determines what she can communicate is an assumption of the model, discussed in detail in Section 2.1.
For the purposes of this example, suppose that there are only two possible messages that the sender can use: call them $m_{\emptyset}$ and $m_{G}$. Assume that $m_{\emptyset}\in M(s)$ for all $s\in[0,1]$, so that it can be interpreted as a ‘silent’ message, meaning that the sender can always choose to say nothing to the receiver. Assume instead that $m_{G}\in M(s)$ iff $s\in[0.5,1]$, meaning that $m_{G}$ proves to the receiver that the state has at least a 50% chance of being $1$.
Assume that the sender is biased in favor of state $1$ in the sense that she (weakly) benefits from inducing higher beliefs about the state in the receiver. Specifically, denote by $v(p)$ the payoff the sender obtains from inducing belief $p\in[0,1]$ in the receiver and suppose that $v(p)=0$ for $p<0.4$, $v(p)=1$ for $p\in[0.4,0.8)$ and $v(p)=3$ for $p\geq 0.8$.
The payoff $v$ is a reduced-form representation of the downstream interaction between the sender and the receiver. [^3] In this example, a standard microfoundation involves the receiver facing a decision problem under uncertainty with three actions: $a_{L}$, $a_{M}$ and $a_{H}$. Her payoff is such that she optimally chooses $a_{L}$ at beliefs strictly below $0.4$, $a_{M}$ at beliefs in $[0.4,0.8)$ and $a_{H}$ at beliefs at or above $0.8$. The sender has state-independent preferences over the receiver’s actions, obtaining payoffs of 0, 11 and $3$ from actions $a_{L}$, $a_{M}$ and $a_{H}$, respectively. The latter’s behavior is not modeled explicitly: the receiver is an active player only insofar as we will require her posterior belief after every message to be always consistent with the verifiability structure (so, after observing $m_{G}$, she must believe the state has at least a $50\%$ chance of being equal to $1$) and formed using Bayes’ rule on the equilibrium path.
Equilibrium unraveling and skepticism.
Suppose first that the sender chooses to acquire information about the state through a signal ‘splitting’ her prior into posteriors 0 and $0.8$. Note that this would be the equilibrium signal if the sender’s information acquisition were public—as in Kamenica and Gentzkow (2011)—attaining an expected payoff of $(\operatorname*{cav}v)(1/4)$, as depicted in Figure 1(a).
[…]
(a) Bayesian persuasion.
[…]
(b) Covert learning and disclosure.
Figure 1: On the left: the sender payoff function $v$ (solid gray) and its concave envelope (thick, dotted) as a function of the receiver’s belief. On the right: verifiability structure $M$ (below the horizontal axis), the sender’s skepticism-adjusted value function (solid black)—overlaid over $v$ (solid gray)—and its concave envelope (thick, dotted). The large black dot corresponds to the sender’s equilibrium expected payoff in each case.
This cannot be an equilibrium signal in my model. To see why, suppose indeed the sender chose such a signal in equilibrium, sending message $m_{\emptyset}$ when her type is 0 (i.e., when she privately discovers that the state is 0 for sure) and message $m_{G}$ when her type is $0.8$ (i.e., when the signal realization is $s=0.8$). In this candidate equilibrium, the sender would obtain a payoff of 0 following message $m_{\emptyset}$, and a payoff of $3$ following message $m_{G}$.
However, the sender could profitably deviate to a signal splitting the prior between 0 and $0.8-\delta$ (for $\delta<0.3$); in the former case sending message $m_{\emptyset}$ and in the latter sending message $m_{G}$. This deviation is not detectable by the receiver (since she does not observe the sender’s choice of signal) and it increases the probability of sending $m_{G}$ while decreasing the probability of sending $m_{\emptyset}$, thus strictly increasing the sender’s expected payoff. This kind of deviation—which worsens the meaning of message $m_{G}$—plays a significant role in shaping equilibrium behavior.
Indeed, only a signal splitting the prior into posteriors 0 and $0.5$ is not prone to such a deviation. These are the lowest sender types consistent with messages $m_{\emptyset}$ and $m_{G}$, respectively, and therefore these messages’ meaning cannot be worsened any further. This signal is part of what I call an ‘unraveling’ equilibrium, in which sender type 0 sends message $m_{\emptyset}$, and sender type $0.5$ sends $m_{G}$. Anticipating this behavior by the sender, the receiver is maximally skeptical, assigning to each message the lowest sender belief that is consistent with it: belief 0 following message $m_{\emptyset}$ and belief $0.5$ following message $m_{G}$. These correspond precisely to the sender’s beliefs, meaning that all acquired information is revealed to the receiver. This equilibrium is depicted in Figure 1(b).
The three features of the unraveling equilibrium just described (on-path sender types being ‘lowest-consistent’ with the messages they send, the receiver’s maximal skepticism and full revelation of acquired information) generalize beyond the specific payoff function $v$ and verifiability structure $M$ of the example. The first result of the paper (Theorem 1) shows that, when $v$ is non-decreasing, all equilibria involving some variation in the receiver’s belief which is payoff-relevant for the sender are ‘unraveling’ in the sense that they satisfy these three properties.
A direct consequence is that the sender’s equilibrium expected payoff is the same across all unraveling equilibria and can be obtained as the concave envelope of a skepticism-adjusted value of posteriors. This function is depicted in Figure 1(b) in the context of the example, and this observation is generalized in Corollary 1.
Uniqueness of unraveling.
In the example, the unraveling equilibrium just presented is, in fact, the only equilibrium. Theorem 2 provides a mild sufficient condition on the primitives of the environment ensuring that all equilibria are unraveling and that therefore the sender’s equilibrium expected payoff is unique.
@@ 1.1 Example and overview of the results
The condition is that the sender must be able to ‘prove news better than the prior’ (PNBP). This property holds if there exists a message that, even when interpreted most skeptically by the receiver, leads to a payoff for the sender strictly above what she would obtain at the prior. In the example, PNBP holds because message $m_{G}$—even if followed by a receiver belief of 0.50.5—leads to a payoff of $1$ for the sender, which is strictly higher than her payoff of 0 at the prior.
At a high level, PNBP ensures that all equilibria are unraveling because a message guaranteeing the sender a payoff strictly above the prior always opens the door to a deviation that worsens the meaning of some message, unless the receiver is maximally skeptical on-path.
@@
Theorem 2 also shows that, in contrast, if the sender cannot prove news better than the prior, there is always a sender-preferred equilibrium in which she acquires and transmits no information. This demonstrates how her bias makes information control worthless, absent some minimal amount of verifiability, as captured by PNBP.
More verifiability and more commitment power.
A natural question to ask is whether (and in which sense) the sender benefits from being able to ‘prove more’. For concreteness, suppose that in the example the sender has access to a new message, call it $m_{E}$, which she can only send if she discovers the excellent news that the state has at least a 90% chance of being high (i.e., $m_{E}\in M(s)$ iff $s\in[0.9,1]$).
It can be easily verified that in equilibrium, the sender will now choose to split the prior between 0 and $0.9$, sending messages $m_{\emptyset}$ and $m_{E}$, respectively. This leads to an increase in the sender’s equilibrium expected payoff compared to the case in which only $m_{\emptyset}$ and $m_{G}$ were available, as depicted in Figure 2.
[…]
Figure 2: Adding message $m_{E}$ makes the sender better off.
@@
What is the appropriate notion of ‘more verifiability’ which ensures that the sender’s payoff improves in equilibrium more generally? Theorem 3 provides a pre-order on the set of verifiability structures that characterizes those shifts that are beneficial for the sender when she can prove news better than the prior. The necessary and sufficient condition is easy to describe: the set of sender types that are able to separate from all lower types with a single message, regardless of the receiver’s belief, must grow in the sense of set inclusion. This guarantees that any equilibrium distribution of receiver beliefs inducible in the lower-ranked verifiability structure is also inducible in the higher-ranked one.
If the sender cannot prove news better than the prior—meaning that unraveling equilibria need not exist—I show instead that shifts towards more verifiability may in fact hurt the sender ex ante (Section 4.3). This is caused by the lack of observability of the sender’s information acquisition strategy. In a more constrained environment, the equilibrium meaning of some messages may worsen with more verifiability: if the sender is known to be able to prove good news, any claim of ignorance she makes will no longer be taken at face value by the receiver.
@@
The comparative statics result is useful for two reasons. First, the pre-order makes it possible to compare a rich variety of common verifiability environments. For example, according to the pre-order, cheap talk is minimal, while full-verifiability structures are maximal. Interval-partition verifiability structures (Example 4)—capturing settings with ‘grades’ or ‘ratings’ that certifiably pool together sets of types—are ranked if and only if one is a refinement of the other (Example 7 and Remark 2). Adding a new message as in the example—thereby allowing the sender to credibly communicate her acquired information more finely—always leads to a higher-ranked structure (Example 8 and Remark 2).
Second, the result can be viewed as providing a tool for comparing Bayesian persuasion environments with different forms of sender commitment. In this context, it characterizes the appropriate notion of ‘increasing commitment power’ which is valuable for the sender.
Optimal verifiability.
@@
The comparative statics result also leads directly to a characterization of sender-optimal verifiability structures (Proposition 1), which allow the sender to attain the full-commitment value in equilibrium. The result therefore highlights how the full commitment assumption in the canonical Bayesian persuasion setting can be relaxed, without affecting the sender’s equilibrium payoff. The necessary and sufficient condition states that all sender types must be able to separate from all lower ones by means of a single message, no matter what the receiver believes. Given the monotone bias of the sender, this guarantees that separation is possible for all types in equilibrium, allowing the sender to replicate the Bayesian persuasion outcome.
@@
Finally, I characterize structures that lead to full information provision in equilibrium (Proposition 2). These are receiver-optimal, provided that the receiver values information about the state. They take a simple form: only the very best news (that is: claiming that the state is equal to $1$ with certainty) is considered credible, while all inconclusive news is fully unverifiable. Every structure of this kind incentivizes perfect information acquisition by the sender, which she then fully reveals in equilibrium.
@@ 1.2 Literature
There is an extensive literature studying the role of verifiability (or lack thereof) in sender-receiver games. A number of papers have pointed out that the classical unraveling result (Viscusi (1978), Grossman and Hart (1980), Grossman (1981) and Milgrom (1981)) breaks down when the sender’s private information is not fully verifiable, in the sense that some sender types do not have access to messages that allow them to sustain separation in equilibrium. [^4] This approach traces back to Dye (1985) and Jung and Kwon (1988) in sender-receiver games and to Green and Laffont (1986) in a mechanism design setting. Important contributions include Okuno-Fujiwara et al. (1990), Lipman and Seppi (1995) and, more recently, Mathis (2008), Hagenbach et al. (2014) and Rappoport (2024). Complete lack of verifiability is studied in the literature on cheap talk, initiated by Crawford and Sobel (1982). In this literature, the sender’s type distribution is fixed and common knowledge; in my model, the players start equally uninformed and the sender can privately acquire information before communicating with the receiver. This paper therefore explores the role of (partially) verifiable information in a context in which what the sender privately knows is a covert strategic decision on her part.
@@
In doing so, this paper contributes to a growing literature which augments sender-receiver games by endogenizing the sender’s private information. This includes papers studying sender learning in cheap talk models, [^5] See, for example, Pei (2015) and Deimen and Szalay (2019). Earlier, Austen-Smith (1994) explored this question in a setting in which the sender either learns the value of the state perfectly, or not at all; in his setting, information acquisition is covert, but the sender can prove that she is informed. settings with full verifiability [^6] An early example is Matthews and Postlewaite (1985), in which the sender faces a binary choice between a fully informative and verifiable signal and a fully uninformative and unverifiable signal. More recently, Gentzkow and Kamenica (2017) show that under full verifiability the sender perfectly reveals all acquired information. and gradual evidence acquisition. [^7] For example, Argenziano et al. (2016), Felgenhauer and Loerke (2017), Janssen (2018), Herresthal (2022). This paper studies the case in which the sender’s covertly acquired private information is only partially verifiable. For this reason, it is closest to a pair of papers studying covert information acquisition in Dye (1985)-like environments: Ben-Porath et al. (2018) and, in particular, DeMarzo et al. (2019). [^8] Information acquisition with Dye evidence is also studied in Kartik et al. (2017) and Shishkin (2024).
As in these papers, my model studies how the lack of observability of the sender’s information acquisition determines what she chooses to learn and communicate with the receiver. There are two substantial differences in our approaches.
The first concerns the class of verifiability structures studied. In the Dye setting these papers consider, the only unverifiable message is ‘non-disclosure’. Its meaning must therefore be determined in equilibrium and will be shaped by the sender’s ability to conceal negative evidence and to covertly change the way she decides to acquire information. In my model, the meaning of several messages—potentially of all of them—must be jointly determined in equilibrium. [^9] More formally, the receiver’s belief following any $m$ for which $M^{-1}(m)$ is not a singleton is determined in equilibrium. With Dye evidence, only the ‘non-disclosure’ message, call it $m_{\emptyset}$, is such that $M^{-1}(m_{\emptyset})$ contains more than one element. My approach to partial verifiability allows me to capture Dye-style evidence (see Example 5) but also to speak directly to a variety of canonical settings in the communication literature which Dye evidence cannot capture. For example, my formalization of verifiability encompasses settings ranging from mandatory disclosure (and, more broadly, full verifiability) to cheap talk (see Examples 1, 2, 3 and 4).
The second key difference is that in these papers the sender’s payoff is affine in the receiver’s belief. It follows that the ‘selective information acquisition’ persuasion motive present in my model (and central to the Bayesian persuasion literature) is absent from these papers.
One important dimension in which my setting is less general than theirs is that they study problems in which the sender is constrained in the signals she can choose. In my model the sender is fully unrestricted in her information acquisition strategy. I discuss this in further detail in the context of Example 5 on 1.
@@
This paper also contributes to the literature exploring the consequences of relaxing the commitment assumption in Kamenica and Gentzkow (2011). Some papers allow the sender to alter negative results with some positive probability (for example, Lipnowski et al. (2022), Fréchette et al. (2022) and Min (2021)) while others study models of Bayesian persuasion with lying costs. In this latter category, Nguyen and Tan (2021) analyze a setting in which the sender’s information acquisition strategy is observable by the receiver while in Guo and Shmaya (2021)’s set-up, as in mine, it is not. Both provide conditions on the costs such that the sender-preferred equilibrium exhibits full information revelation.
@@
In contrast to these papers, in my model the sender faces a constraint on what she can or cannot communicate depending on what she discovers. This distinct formalization of partial commitment allows me to bridge the gap between the Bayesian persuasion setting and models of verifiable disclosure. This, in turn, leads to a characterization of when ‘more verifiability’ of the sender’s acquired information is akin to her having ‘more commitment power’.
@@
I introduce the model in the next section. Unpersuasive and unraveling equilibria are studied in Section 3, along with a discussion of the value of persuasion in the context of my model of disclosure with covert learning. Section 4 contains the comparative statics result, characterizing changes in verifiability that are beneficial for the sender, and the characterization of sender-optimal verifiability structures. Receiver-optimal structures are characterized in Section 5.